OpenStudy (anonymous):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (anonymous):

put it in y = mx+b form in other words solve the equation for y so you have y on the left of the equal sign and everything else on the right side.

OpenStudy (anonymous):

slope = -3/5 ; y-intercept = (0, -6)

OpenStudy (anonymous):

5x - 3y = -6 3y = 5x+6 y = 5/3x + 2 slope: 5/3 y intercept: (0,2)

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

one more?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Write the equation of the line in slope-intercept form that passes through the point (3, -2) and has a slope of 2